# Order isomorphisms of sup-stable function spaces: continuous, Lipschitz, c-convex, and beyond

### v2

@@ 3 Studying (max,+) and order isomorphisms over specific sets
Applying the results of Section 2, and in particular 21, we now characterize the (max,+)-isomorphisms for some given sets of functions: i) lower semicontinuous or continuous, ii) 1-Lipschitz with a (weak) locally compact or complete metric, iii) lower semicontinuous and convex. For some of these sets we can even characterize the order isomorphisms in great generality.
@@ 3.1 Indicator functions and lower semicontinuous functions
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, for $\mathscr{X}$ compact, by 15, since its sup-closure is the set of proper lower semicontinuous functions, the (max,+)-isomorphisms between spaces of continuous functions are of the form (27). The converse part is immediate.∎
@@
Concerning lattice isomorphisms $J$ of continuous functions over compact Hausdorff spaces, assuming $J$ to be continuous under pointwise convergence, which is weaker than (13), formula (28) was first shown in [Kap48, Theorem 1] with an elegant proof based on characterizing “prime ideals” of $C(\mathscr{X},\mathbb{R})$. When the continuity assumption is dropped, formula (28) still holds on a $G_{\delta}$-dense set [CS08, Theorem 1, Lemma 4], which encompasses the results of Kaplansky. In [LT16, Theorem 4.5], the same expression was obtained for isomorphims of some specific vector subspaces of continuous functions on a metric space, named “near vector lattices” by the authors, and encompassing (locally) Lipschitz or uniformly continuous (bounded) functions. Compactifications of $\mathscr{X}$ and continuity of $J$, which is not automatic, are key elements in this approach. They do not appear when focusing on lower semicontinuous functions.
@@
The $\mathcal{G}$-relatively-inf-irreducible points of lower semicontinuous functions are the Dirac masses $(\delta_{x}+\lambda)_{x\in\mathscr{X},\lambda\in\overline{\mathbb{R}}}$. Indeed if $g,h\in\overline{\mathbb{R}}^{\mathscr{Y}}$ are such that $\delta_{x}=\inf\nolimits^{\mathcal{G}}(g,h)$, then $\delta_{x}\leq g$ and $\delta_{x}\leq h$, so both are Dirac masses $\delta_{x}+\lambda$, one being equal to $\delta_{x}$. Conversely take a $\mathcal{G}$-relatively-inf-irreducible $f\in\mathcal{G}$. It is necessarily inf-irreducible by 5. Assume there exists $x_{0}\neq x_{1}$ such that $f(x_{0})<\infty$ and $f(x_{1})<\infty$, then fix a closed subset $F\subset\mathscr{X}$ such that $x_{0}\in F$, $x_{0}\notin\overline{\mathscr{X}\backslash F}$ and $x_{1}\notin F$. Hence $f=\inf(f+\delta_{F},f+\delta_{\overline{\mathscr{X}\backslash F}})$, both of these functions being different than $f$. Thus $f=\delta_{x}+\lambda$ for some $\lambda\in\overline{\mathbb{R}}$.
Fix $y\in\mathscr{Y}$ and $\lambda\in\mathbb{R}$. Using 13, we have that $J$ maps $\delta_{y}+\lambda$ on some $\delta_{\phi(y,\lambda)}+g_{y}(\lambda)$, for some $\phi:\mathscr{Y}\times\mathbb{R}\rightarrow\mathscr{X}$ and $g:\mathscr{Y}\times\mathbb{R}\rightarrow\mathbb{R}$, the $\infty$-valued functions being mapped onto themselves.
@@
Let $F$ be a closed set of $\mathscr{Y}$, then $\delta_{F}$ is lower semicontinuous, so the sum $g^{1}(\cdot,0)+\delta_{F}(\cdot)$ is also lower semicontinuous. We have that $J(g^{1}(\cdot,0)+\delta_{F}(\cdot))(x)=g_{x}(g^{1}(\phi(x),0))=0$ for $x$ such that $\phi(x)\in F$ and $\infty$ otherwise, since $g_{x}(\infty)=\infty$ by increasingness and bijectivity of $g_{x}(\cdot)$. Hence $J(g^{1}(\cdot,0)+\delta_{F}(\cdot))=\delta_{\phi^{-1}(F)}$ is lower semicontinuous and $\phi^{-1}(F)$ is closed. Consequently $\phi$ is bijective and continuous, repeating the proof, the same holds for $\phi^{-1}$, thus $\phi$ is a homeomorphism.
@@
[…]
@@
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, for $\mathscr{X}$ compact and $J$ satisfying (13), by 15, since its sup-closure is the set of proper lower semicontinuous functions, the order isomorphisms between spaces of continuous functions are of the form (29). In all the proof above, we just replace lower semicontinuous by continuous with no other change to derive the result applying 15.∎
About anti-involutions, if $\mathscr{X}$ is discrete, then $-\operatorname*{Id}$ is an anti-involution over (non proper) lower semicontinuous functions. However if the space has no isolated points then there are no anti-involutions.
@@
The intuition here is that anti-isomorphisms send relative-inf-irreducible points onto sup-irreducible points by 13. However the sup-irreducible points of lower semicontinuous functions are only the $\delta^{\bot}_{x}$ of isolated $x$.
@@
Let us first characterize the sup-irreducible points. Let $f$ be a lower semicontinuous function taking finite values at two distinct points $x_{0}$ and $x_{1}$. Define two functions $\chi_{0},\chi_{1}$ by setting $\chi_{0}(x_{0})=-1$, $\chi_{1}(x_{1})=-1$ and 00 elsewhere. Then $f=\sup(f+\chi_{0},f+\chi_{1})$, with the two latter functions being lower semicontinuous as a sum of such functions, so $f$ is not sup-irreducible. The same applies if $f$ takes the +$\infty$+ ∞ value at two distinct points $x_{0}$ and $x_{1}$ by setting $f_{0}=f$ everywhere except at $x_{0}$ where we choose $f_{0}(x_{0})=0$ (resp. $f_{1}(x_{1})=0$), so $f=\sup(f_{0},f_{1})$. The function $f_{0}$ is l.s.c. since its epigraph is the union of that of $f$ and of $\{x_{0}\}\times[0,+\infty]$. This shows that every sup-irreducible function has to be of the form $\delta^{\bot}_{x}+\lambda$ with $\lambda\in\overline{\mathbb{R}}$. For $\delta^{\bot}_{x}$ to be lower semicontinuous, we need to have $\mathscr{X}\backslash\{x\}$ to be closed, i.e. $x$ to be isolated.
Assume there exists an anti-isomorphism $T$. Similarly to the proof of 30, applying 13, it is easy to show that $T(\delta^{\top}_{y}+\lambda)=\delta^{\bot}_{\phi(y)}+g(y,\lambda)$ with $\phi:\mathscr{Y}\to I_{\mathscr{X}}$ bijective and $g(y,\lambda)\in\mathbb{R}$. In the same fashion, we have a bijective $\psi:\mathscr{X}\to I_{\mathscr{Y}}$. Hence $\psi\circ\phi:\mathscr{Y}\to I_{\mathscr{Y}}$ is injective, and by the Cantor–Schröder–Bernstein theorem, $\mathscr{Y}$ and $I_{\mathscr{Y}}$ are thus in bijection. Considering $T^{-1}$, the sets $\mathscr{X}$ and $I_{\mathscr{X}}$ are also in bijection. So we can construct with these bijections and the previous injections, two injections from $\mathscr{X}$ to $\mathscr{Y}$ and $\mathscr{Y}$ to $\mathscr{X}$. Again by the Cantor–Schröder–Bernstein theorem, we derive a bijection between $\mathscr{X}$ and $\mathscr{Y}$.
Note that 32 does not rule out that there can be interesting anti-isomorphisms $T:\mathscr{F}\to\mathcal{G}$ over lower semicontinuous functions $\mathscr{F}$ with $\mathcal{G}$ another function space, as given for instance by [Bac01, Theorem 2.2] for $\mathscr{X}$ a Banach space.
@@ 3.2 Busemann points and 1-Lipschitz functions
In [CSCS11, Theorem 1], it was first shown that, for Lipschitz functions on bounded metric spaces, the structure of order isomorphisms was necessarily that of lower semicontinuous functions, obtained in 30, with a bi-Lipschitz $\phi$. Extending this result to general metric spaces, [LT16, Theorem 5.5, eq.(8)] provided an “if an only if” condition on the existence of a bi-Lipschitz $\phi$. However at the moment, to the best of our knowledge there is no sufficient condition on the form of the function $g:\mathscr{X}\times\mathbb{R}\rightarrow\mathbb{R}$, not even lower semicontinuity, [^5] On Lipschitz functions, even for $\phi=\operatorname*{Id}$, the monograph [AZ90, Chapter 7] says: “Quite amazing is the fact that the operator $J$ may act in a Hölder space, although the generating function $g$ is not continuous.” Further conditions, like global Lipschitzianity of $J$, can however ensure that $J$ is affine as in [AZ90, Theorem 7.8] for $\mathscr{X}=[0,1]$. The necessity of extra assumptions on $J$ is also discussed therein if one is interested by differentiability properties. nor any specific results for order isomorphisms of the space of $1$-Lipschitz functions.
Nevertheless a similar result to what follows was shown in [CSJ17, Corollary 4.1] for 1-Lipschitz functions over complete metric space, for order isomorphisms that also preserve convex combinations, i.e. $J(sf+(1-s)f^{\prime})=sJf+(1-s)Jf^{\prime}$, through very different proof techniques such as bases of topologies. The line of research of Cabello et al. on weak metrics was adapted later in [DJVM20, Theorem 3.1] to Finsler manifolds. We instead do not ask that convexity is preserved but that $J$ commutes with the addition of constants.
@@
Unlike [CSJ17, DJVM20], we do not assume the weak metric to be nonnegative-valued, which allows us to consider the Funk metric in 40, for which $\delta_{s}$ is the Hilbert metric. We still assume that $\delta_{s}$ is a metric, hence has nonnegative values.
@@
For $p\in(0,1]$ and $\delta(x,y):\mathscr{X}\times\mathscr{X}\to[0,+\infty)$, by subadditivity of $t\in\mathbb{R}_{+}\mapsto t^{p}$, the result above also applies to $p$-Hölder functions since $\delta^{p}$ is also a weak metric. Other interesting examples of weak metrics are those of the form “$\inf_{\gamma}\int L(\gamma)$” with nonnegative Lagrangian $L$, used in [Vil09, Chapter 7] or [AFG24, Section 7.2].
The proof of 34 uses the compactification of $(\mathscr{X},\delta)$ by horofunctions and is delayed to p.3.2, after showing some intermediary lemmas. The horofunction compactification was introduced by Gromov [Gro81], see also [BGS85], [Bal95, Ch. II] and [Rie02]. It carries over to weak metrics, see [AGW09, GV12, Wal18]. For compact $\mathscr{X}$, 34 is merely a specialization of 27.
@@
Notice that $-\operatorname*{Id}$ is an anti-involution over $1$-Lipschitz functions, and just apply 11 in conjunction with (30). ∎
@@
By the triangle inequality, $-\delta(y,\bar{x})\leq\imath(x)(y)\leq\delta(y,\bar{x})$, so that $i(\mathscr{X})$ can be identified to a subset of the Cartesian product $\prod_{y\in\mathscr{X}}[-\delta(y,\bar{x}),\delta(y,\bar{x})]$ which is compact in the product topology, by Tychonoff’s theorem. The horocompactification of $\mathscr{X}$ is defined as the closure $\overline{\imath(\mathscr{X})}$ of $\imath(\mathscr{X})$ in the product topology. As a special case of a “second” Ascoli’s theorem [Sch70, p310], since $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X},\mathbb{R})$ with the pointwise convergence topology is isomorphic to itself with the compact-open topology and thus with uniform convergence on compact sets, the closure of $\imath(\mathscr{X})$ is the same if we take the topology of uniform convergence on compact sets of the metric space $(\mathscr{X},\delta_{s})$. We set $\mathscr{X}(\infty)\coloneqq\overline{\imath(\mathscr{X})}\setminus\imath(\mathscr{X})$. An element of $\mathscr{X}(\infty)$ is called a horofunction.
@@
As already discussed in 24, inf-stable complete subspaces of ¯$\overline{\mathbb{R}}^{\mathscr{X}}$ are sets of non-expansive functions for the $\overline{\mathbb{R}}$-valued weak metric given by $e_{x}$, so 37 can be easily adapted to this abstract setting.
@@
$-\delta_{s}(x,x_{\alpha})\leq\delta(y,x_{\alpha})-\delta(\bar{x},x_{\alpha})-(\delta(y,x)-\delta(\bar{x},x))\leq\delta_{s}(x,x_{\alpha}).$
@@
Case (ii). Assume that $\delta_{s}(\bar{x},x_{\alpha})$ has a bounded subnet, w.l.o.g. we restrict to this subnet. Take $\alpha\leq\alpha^{\prime}$, then using the triangular inequality and the almost geodesic property, for any $\epsilon>0$ and for $\alpha$ large enough
@@
Since $\delta=\frac{1}{2}\delta_{s}$ and $(X,\delta_{s})$ is complete, take any $\alpha,\alpha^{\prime}$ large enough, and set $\beta=\max(\alpha,\alpha^{\prime})$, then
@@
By 37, $h$ is the limit of $\imath(x_{\alpha})$ where $x_{\alpha}$ is an almost-geodesic net such that $\delta_{s}(\bar{x},x_{\alpha})$ tends to infinity. Let us fix $\epsilon>0$, so that for all $\alpha,\beta$ large enough
@@
and so $a-1+\delta(x_{\alpha},\bar{x})+\delta(\bar{x},x_{\alpha})=a-1+\delta_{s}(\bar{x},x_{\alpha})\leq\epsilon$. However, by 37, the almost-geodesic $x_{\alpha}$ is such that $\delta_{s}(\bar{x},x_{\alpha})$ is unbounded, so we arrive at a contradiction.∎
@@
The compactness or completeness assumption is useful in 38. Indeed, if $\mathscr{X}=\mathbb{Q}$, then, every element $z\in\mathbb{R}\setminus\mathbb{Q}$ yields an inf-irreducible horofunction $(d(\cdot,z)-d(\bar{x},z))\in\operatorname*{\operatorname{Lip}_{1}}(\mathbb{Q})$ which is in the same Archimedean class than every map $\imath(x)$ with $x\in\mathscr{X}$, while having $d(\cdot,z)\notin\imath(\mathscr{X})$. 38 allows to separate the Busemann points from the distance functions $\imath(\mathscr{X})$ based on the domination criterion $h\geq a+\delta(\cdot,\bar{x})$. For instance for the space of 1-Lipschitz functions over a Banach space $\mathscr{X}$, the Busemann points are the inf-irreducible $\langle p,\cdot\rangle$ for $\|p\|_{\mathscr{X}^{*}}=1$, and we can thus exclusively focus on the inf-irreducible functions $e_{x}=\delta(\cdot,x)$ with maximal Archimedean class.
Proof of 34.
A (max,+) isomorphism $J$ of $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X})$ must preserve inf-irreducible elements. Hence, by 36, denoting by $\mathcal{B}(\mathscr{X})$ the set of Busemann points of $\mathscr{X}$, $J$ must preserve the set $\mathbb{R}+(\imath(\mathscr{X})\cup\mathcal{B}(\mathscr{X}))$. Owing to the triangular inequality, the elements of $\imath(\mathscr{X})$ all belong to the same Archimedean class, which is maximal. By 16, a (max,+) isomorphism must send a maximal Archimedean class to a maximal Archimedean class. By 38, the Archimedean class of a Busemann point is not maximal. We conclude that $J$ preserves the set $\mathbb{R}+\imath(\mathscr{X})$.
By 21, we thus have (23), that is $Jf=g+f\circ\phi$. We just have to characterize the properties of $g$ and $\phi$. Since $J$ sends nonexpansive maps onto nonexpansive maps, we have that, for all nonexpansive maps $f:\mathscr{Y}\rightarrow\mathbb{R}$, and all $x,x^{\prime}\in\mathscr{X}$
@@
Setting $f=\delta^{\prime}(\cdot,\phi(x^{\prime}))$, since $02\delta^{\prime}(\phi(x),\phi(x))=\delta^{\prime}(\phi(x),\phi(x))=$ 0, we get that for all $x,x^{\prime}\in\mathscr{X}$
@@
$-g(\phi^{-1}(y))+g(\phi^{-1}(y^{\prime}))+\delta(\phi^{-1}(y),\phi^{-1}(y^{\prime}))\leq\delta^{\prime}(y,y^{\prime}).$ (34)
@@
$\delta(x,x^{\prime})\leq g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))\leq\delta(x,x^{\prime}).$ (35)
@@
Let $C$ be a closed convex pointed reproducing cone in $\mathbb{R}^{n}$. Such a cone defines a partial order $\leq$ on $\mathbb{R}^{n}$ such that $x\leq y$ iff $y-x\in C$. We define the (reverse) Funk “metric” in the interior of $C$ as follows:

### v3

@@ 3 Studying (max,+) and order isomorphisms over specific sets
Applying the results of Section˜2, and in particular ˜21, we now characterize the (max,+)-isomorphisms for some given sets of functions: i) lower semicontinuous or continuous, ii) 1-Lipschitz with a (weak) locally compact or complete metric, iii) lower semicontinuous and convex. For some of these sets we can even characterize the order isomorphisms in great generality.
@@ 3.1 Indicator functions and lower semicontinuous functions
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, for $\mathscr{X}$ compact, by ˜15, since its sup-closure is the set of proper lower semicontinuous functions, the (max,+)-isomorphisms between spaces of continuous functions are of the form (26). The converse part is immediate.∎
@@
Concerning lattice isomorphisms $J$ of continuous functions over compact Hausdorff spaces, assuming $J$ to be continuous under pointwise convergence, which is weaker than (13), formula (27) was first shown in [Kap48, Theorem 1] with an elegant proof based on characterizing “prime ideals” of $C(\mathscr{X},\mathbb{R})$. When the continuity assumption is dropped, formula (27) still holds on a $G_{\delta}$-dense set [CS08, Theorem 1, Lemma 4], which encompasses the results of Kaplansky. In [LT16, Theorem 4.5], the same expression was obtained for isomorphisms of some specific vector subspaces of continuous functions on a metric space, named “near vector lattices” by the authors, and encompassing (locally) Lipschitz or uniformly continuous (bounded) functions. Compactifications of $\mathscr{X}$ and continuity of $J$, which is not automatic, are key elements in these approaches. They do not appear when focusing on lower semicontinuous functions.
@@
[…]
Fix $y\in\mathscr{Y}$ and $\lambda\in\mathbb{R}$. Using ˜13, we have that $J$ maps $\delta_{y}+\lambda$ on some $\delta_{\phi(y,\lambda)}+g_{y}(\lambda)$, for some $\phi:\mathscr{Y}\times\mathbb{R}\rightarrow\mathscr{X}$ and $g:\mathscr{Y}\times\mathbb{R}\rightarrow\mathbb{R}$, the $\infty$-valued functions being mapped onto themselves.
@@
Let $F$ be a closed subset of $\mathscr{Y}$, then $\delta_{F}$ is lower semicontinuous, so the sum $g^{1}(\cdot,0)+\delta_{F}(\cdot)$ is also lower semicontinuous. We have that $J(g^{1}(\cdot,0)+\delta_{F}(\cdot))(x)=g_{x}(g^{1}(\phi(x),0))=0$ for $x$ such that $\phi(x)\in F$ and $\infty$ otherwise, since $g_{x}(\infty)=\infty$ by increasingness and bijectivity of $g_{x}(\cdot)$. Hence $J(g^{1}(\cdot,0)+\delta_{F}(\cdot))=\delta_{\phi^{-1}(F)}$ is lower semicontinuous and $\phi^{-1}(F)$ is closed. Consequently $\phi$ is bijective and continuous, repeating the proof, the same holds for $\phi^{-1}$, thus $\phi$ is a homeomorphism.
@@
$=\liminf_{n\rightarrow\infty}\liminf_{m\rightarrow\infty}g(x_{n},f(x_{m}))\leq\liminf_{n\rightarrow\infty}g(x_{n},f(x_{n})).$
@@
Concerning the set of continuous functions $\mathcal{C}(\mathscr{X},\mathbb{R})$, for $\mathscr{X}$ compact and $J$ satisfying (13), by ˜15, since its sup-closure is the set of proper lower semicontinuous functions, the order isomorphisms between spaces of continuous functions are of the form (28). In all the proof above, we just replace lower semicontinuous by continuous with no other change to derive the result applying ˜15.∎
About anti-involutions, if $\mathscr{X}$ is discrete, then $\operatorname*{Id}$ is an anti-involution over (non proper) lower semicontinuous functions. However if the space has no isolated points then there are no anti-involutions.
@@
The intuition here is that anti-isomorphisms send relative-inf-irreducible points onto sup-irreducible points by ˜13. However the sup-irreducible points of lower semicontinuous functions are only the $\delta^{\bot}_{x}$ of isolated $x$.
@@
[…]
Assume there exists an anti-isomorphism $T$. Similarly to the proof of ˜29, applying ˜13, it is easy to show that $T(\delta^{\top}_{y}+\lambda)=\delta^{\bot}_{\phi(y)}+g(y,\lambda)$ with $\phi:\mathscr{Y}\to I_{\mathscr{X}}$ bijective and $g(y,\lambda)\in\mathbb{R}$. In the same fashion, we have a bijective $\psi:\mathscr{X}\to I_{\mathscr{Y}}$. Hence $\psi\circ\phi:\mathscr{Y}\to I_{\mathscr{Y}}$ is injective, and by the Cantor–Schröder–Bernstein theorem, $\mathscr{Y}$ and $I_{\mathscr{Y}}$ are thus in bijection. Considering $T^{-1}$, the sets $\mathscr{X}$ and $I_{\mathscr{X}}$ are also in bijection. So we can construct with these bijections and the previous injections, two injections from $\mathscr{X}$ to $\mathscr{Y}$ and $\mathscr{Y}$ to $\mathscr{X}$. Again by the Cantor–Schröder–Bernstein theorem, we derive a bijection between $\mathscr{X}$ and $\mathscr{Y}$.
Note that ˜31 does not rule out that there can be interesting anti-isomorphisms $T:\mathscr{F}\to\mathcal{G}$ over lower semicontinuous functions $\mathscr{F}$ with $\mathcal{G}$ another function space, as given for instance by [Bac01, Theorem 2.2] for $\mathscr{X}$ a Banach space.
@@ 3.2 Busemann points and 1-Lipschitz functions
In [CSCS11, Theorem 1], it was first shown that, for Lipschitz functions on bounded metric spaces, the structure of order isomorphisms was necessarily that of lower semicontinuous functions, obtained in ˜29, with a bi-Lipschitz $\phi$. Extending this result to general metric spaces, [LT16, Theorem 5.5, eq.(8)] provided an “if an only if” condition on the existence of a bi-Lipschitz $\phi$. However at the moment, to the best of our knowledge there is no sufficient condition on the form of the function $g:\mathscr{X}\times\mathbb{R}\rightarrow\mathbb{R}$, not even lower semicontinuity, [^5] On Lipschitz functions, even for $\phi=\operatorname*{Id}$, the monograph [AZ90, Chapter 7] says: “Quite amazing is the fact that the operator $J$ may act in a Hölder space, although the generating function $g$ is not continuous.” Further conditions, like global Lipschitzianity of $J$, can however ensure that $J$ is affine as in [AZ90, Theorem 7.8] for $\mathscr{X}=[0,1]$. The necessity of extra assumptions on $J$ is also discussed therein if one is interested by differentiability properties. nor any specific results for order isomorphisms of the space of $1$-Lipschitz functions.
Nevertheless a similar result to what follows was shown in [CSJ17, Corollary 4.1] for 1-Lipschitz functions over complete metric spaces, for order isomorphisms that also preserve convex combinations, i.e. $J(sf+(1-s)f^{\prime})=sJf+(1-s)Jf^{\prime}$, through very different proof techniques such as bases of topologies. The line of research of Cabello et al. on weak metrics was adapted later in [DJVM20, Theorem 3.1] to Finsler manifolds. We instead do not ask that convexity is preserved but that $J$ commutes with the addition of constants.
@@
Unlike [CSJ17, DJVM20], we do not assume the weak metric to be nonnegative-valued, which allows us to consider the Funk metric in ˜39, for which $\delta_{s}$ is the Hilbert metric. We still assume that $\delta_{s}$ is a metric, hence has nonnegative values. Our result reads as follows
@@
For $p\in(0,1]$ and $\delta(x,y):\mathscr{X}\times\mathscr{X}\to[0,+\infty)$, by subadditivity of $t\in\mathbb{R}_{+}\mapsto t^{p}$, the result above also applies to $p$-Hölder functions since $\delta^{p}$ is also a weak metric. Other interesting examples of weak metrics are those of the form $\inf_{\gamma}\int L(\gamma)$” with nonnegative Lagrangian $L$, used in [Vil09, Chapter 7] or [AFG24, Section 7.2].
The proof of ˜33 uses the compactification of $(\mathscr{X},\delta)$ by horofunctions and is delayed to p.3.2, after showing some intermediary lemmas. The horofunction compactification was introduced by Gromov [Gro81], see also [BGS85], [Bal95, Ch. II] and [Rie02]. It carries over to weak metrics, see [AGW09, GV12, Wal18]. For compact $\mathscr{X}$, ˜33 is merely a specialization of ˜27.
@@
Notice that $\operatorname*{Id}$ is an anti-involution over 11-Lipschitz functions, and just apply ˜11 in conjunction with (29). ∎
@@
By the triangle inequality, $\delta(y,\bar{x})\leq\imath(x)(y)\leq\delta(y,\bar{x})$, so that $i(\mathscr{X})$ can be identified to a subset of the Cartesian product $\prod_{y\in\mathscr{X}}[-\delta(y,\bar{x}),\delta(y,\bar{x})]$ which is compact in the product topology, by Tychonoff’s theorem. The horocompactification of $\mathscr{X}$ is defined as the closure $\overline{\imath(\mathscr{X})}$ of $\imath(\mathscr{X})$ in the product topology. As a special case of a “second” Ascoli’s theorem [Sch70, p310], since $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X},\mathbb{R})$ with the pointwise convergence topology is isomorphic to itself with the compact-open topology and thus with uniform convergence on compact sets, the closure of $\imath(\mathscr{X})$ is the same if we take the topology of uniform convergence on compact sets of the metric space $(\mathscr{X},\delta_{s})$. We set $\mathscr{X}(\infty)\coloneqq\overline{\imath(\mathscr{X})}\setminus\imath(\mathscr{X})$. An element of $\mathscr{X}(\infty)$ is called a horofunction.
@@
As already discussed in Remark 24, inf-stable complete subspaces of $\overline{\mathbb{R}}^{\mathscr{X}}$ are sets of non-expansive functions for the $\overline{\mathbb{R}}$-valued weak metric given by $e_{x}$, so ˜36 can be easily adapted to this abstract setting.
@@
$\delta_{s}(x,x_{\alpha})\leq\delta(y,x_{\alpha})-\delta(\bar{x},x_{\alpha})-(\delta(y,x)-\delta(\bar{x},x))\leq\delta_{s}(x,x_{\alpha}).$
@@
Case (ii). Assume that $\delta_{s}(\bar{x},x_{\alpha})$ has a bounded subnet, w.l.o.g. we restrict to this subnet. Take $\alpha\leq\alpha^{\prime}$, then using the triangle inequality and the almost geodesic property, for any $\epsilon>0$ and for $\alpha$ large enough
@@
So, as $\delta=\frac{1}{2}\delta_{s}$, the almost geodesic property $\delta(\bar{x},x_{\alpha})+\delta(x_{\alpha},x_{\alpha^{\prime}})\leq\epsilon+\delta(\bar{x},x_{\alpha^{\prime}})$ gives that for $\alpha\leq\alpha^{\prime}$ large enough, $\delta(x_{\alpha},x_{\alpha^{\prime}})\leq 2\epsilon$. Using again that $\delta=\frac{1}{2}\delta_{s}$, we have that $\delta_{s}(x_{\alpha},x_{\alpha^{\prime}})\leq 4\epsilon$ for any $\alpha,\alpha^{\prime}$ large enough. Since $(X,\delta_{s})$ is complete, setting $\beta=\max(\alpha,\alpha^{\prime})$, then
@@
By ˜36, $h$ is the limit of $\imath(x_{\alpha})$ where $x_{\alpha}$ is an almost-geodesic net such that $\delta_{s}(\bar{x},x_{\alpha})$ tends to infinity. Let us fix $\epsilon>0$, so that for all $\alpha,\beta$ large enough
@@
and so $a-1+\delta(x_{\alpha},\bar{x})+\delta(\bar{x},x_{\alpha})=a-1+\delta_{s}(\bar{x},x_{\alpha})\leq\epsilon$. However, by ˜36, the almost-geodesic $x_{\alpha}$ is such that $\delta_{s}(\bar{x},x_{\alpha})$ is unbounded, so we arrive at a contradiction.∎
@@
The compactness or completeness assumption cannot be dismissed in ˜37. Indeed, if $\mathscr{X}=\mathbb{Q}$, then, every element $z\in\mathbb{R}\setminus\mathbb{Q}$ yields an inf-irreducible horofunction $(d(\cdot,z)-d(\bar{x},z))\in\operatorname*{\operatorname{Lip}_{1}}(\mathbb{Q})$ which is in the same Archimedean class than every map $\imath(x)$ with $x\in\mathscr{X}$, while having $d(\cdot,z)\notin\imath(\mathscr{X})$. ˜37 allows to separate the Busemann points from the distance functions $\imath(\mathscr{X})$ based on the domination criterion $h\geq a+\delta(\cdot,\bar{x})$. For instance for the space of 1-Lipschitz functions over a Banach space $\mathscr{X}$, the Busemann points are the inf-irreducible $\langle p,\cdot\rangle$ for $\|p\|_{\mathscr{X}^{*}}=1$, and we can thus exclusively focus on the inf-irreducible functions $e_{x}=\delta(\cdot,x)$ with maximal Archimedean class.
Proof of ˜33.
A (max,+) isomorphism $J$ of $\operatorname*{\operatorname{Lip}_{1}}(\mathscr{X})$ must preserve inf-irreducible elements. Hence, by ˜35, denoting by $\mathcal{B}(\mathscr{X})$ the set of Busemann points of $\mathscr{X}$, $J$ must preserve the set $\mathbb{R}+(\imath(\mathscr{X})\cup\mathcal{B}(\mathscr{X}))$. Owing to the triangle inequality, the elements of $\imath(\mathscr{X})$ all belong to the same Archimedean class, which is maximal. By ˜16, a (max,+) isomorphism must send a maximal Archimedean class to a maximal Archimedean class. By ˜37, the Archimedean class of a Busemann point is not maximal. We conclude that $J$ preserves the set $\mathbb{R}+\imath(\mathscr{X})$.
By ˜21, we thus have (23), that is $Jf=g+f\circ\phi$. We just have to characterize the properties of $g$ and $\phi$. Since $J$ sends nonexpansive maps onto nonexpansive maps, we have that, for all nonexpansive maps $f:\mathscr{Y}\rightarrow\mathbb{R}$, and all $x,x^{\prime}\in\mathscr{X}$
@@
Setting $f=\delta^{\prime}(\cdot,\phi(x^{\prime}))$, since $2\delta^{\prime}(\phi(x),\phi(x))=\delta^{\prime}(\phi(x),\phi(x))=0$, we get that for all $x,x^{\prime}\in\mathscr{X}$
@@
$g(\phi^{-1}(y))+g(\phi^{-1}(y^{\prime}))+\delta(\phi^{-1}(y),\phi^{-1}(y^{\prime}))\leq\delta^{\prime}(y,y^{\prime}).$ (33)
@@
$\delta(x,x^{\prime})\leq g(x)-g(x^{\prime})+\delta^{\prime}(\phi(x),\phi(x^{\prime}))\leq\delta(x,x^{\prime})$ (34)
@@
Let $C$ be a closed convex pointed cone in $\mathbb{R}^{n}$. Such a cone defines a partial order $\leq$ on $\mathbb{R}^{n}$ such that $x\leq y$ iff $y-x\in C$. We define the (reverse) Funk “metric” in the interior of $C$ as follows: